OpenStudy (anonymous):

What is the volume of a rectangular sandbox that is 3 feet by 4 feet by 1 foot? 7 ft^ 8 ft^ 12 ft^ 13 ft^

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3 \times 4 \times 1 }\)

jimthompson5910 (jim_thompson5910):

Similar to the last problem, Volume = Length * Width * Height

Parth (parthkohli):

\( \color{Black}{\Rightarrow length \times width \times height }\)

OpenStudy (anonymous):

thanks!

Parth (parthkohli):

yw